determine m such that the two vectors 3i-9j and mi+2j are othogonal
Answer by longjonsilver(2297)

orthogonal means "at right angles". For 2 vectors at right angles, their dot (or scalar) product is zero.
So, (3 -9) dot (m 2) = 0
3m + -18 = 0
3m - 18 = 0
3m = 18
m = 18/3
m = 6